### v1

@@ 5 Conclusion
In this work, we address ensemble-based zero-shot machine translation by directly translating and pivoting through different languages. We further design a novel bi-level beam search algorithm (called EBBS) for decoding. We evaluated EBBS on two popular zero-shot translation datasets, IWSLT and Europarl. Results show that EBBS outperforms existing ensemble techniques, and that the high-quality translations produced by EBBS can be used for distillation to improve translation efficiency (and sometimes also quality over EBBS).
6 Limitations
Our work features both algorithmic design and empirical effectiveness, but may also have limitations.
First, our work focuses on zero-shot translation, due to the background of our project. Nevertheless, the EBBS algorithm that we have developed is a general decoding method and is potentially applicable to various text generation tasks. We are happy to explore this direction as future work, such as building ensembles of large language models.
Second, our distillation experiments were conducted on Italian-to-Dutch translation because it appears in both IWSLT and Europarl datasets. We are unable to perform all-direction distillation in this project, because Europarl, say, has 56 zero-shot directions, and distillation techniques require extensive computing resources to generate pseudo-groundtruth sentences. Nevertheless, our distillation results are consistent on both datasets, and therefore, not including all-direction distillation does not hurt the validity of our method. Moreover, our work points to interesting future work, such as iteratively applying our ensemble-then-distill procedure to build a self-improving multilingual machine translation system.
Third, we have mainly used BLEU (in particular, SacreBLEU) scores to evaluate translation quality, which is standard in multilingual translation research Johnson et al. (2017); Liu et al. (2021); Fan et al. (2021). We are aware of language model-based scores, such as BERTScore Zhang et al. (2019), for evaluating the translation in certain languages, but they may not be appropriate for low-resource languages and are not commonly used in our setting. We have not performed human evaluation either, due to both practical and ethical concerns. Practically, it is difficult to find qualified annotators because our multilingual setting requires an annotator to know a large number of languages in addition to English. Ethically, it is inappropriate to ask annotators to study these languages before or during annotation, which is exhausting and may last for extended periods of time. That being said, our work uses the most standard metrics and results are consistent in most BLEU variants on both datasets, with a strong statistical confidence level based on 62 translation directions. In addition, we have provided case studies to further illustrate how our EBBS works. Therefore, we deem our evaluation ethical, appropriate, and adequate.
Acknowledgment
The research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), a Mitacs Accelerate project, the Amii Fellow Program, the Canada CIFAR AI Chair Program, an Alberta Innovates Program, and the Digital Research Alliance of Canada (alliancecan.ca). In the title, we used the open-domain artwork “The Great Wave off Kanagawa” by Katsushika Hokusai.
@@ 12       $\triangleright$ check for termination
Appendix B Details of Our Experiments
Dataset details. We evaluated our methods using IWSLT 2017 Cettolo et al. (2017) and Europarl v7 Koehn (2005). Table 6 provides a summary of the languages.
The IWSLT 2017 translation dataset features multilingual data derived from TED talks. We followed previous work and used a standard split for zero-shot evaluation (Dabre and Kurohashi, 2017; Liu et al., 2021). In particular, IWSLT contains English-centric training data for Italian, Dutch, and Romanian, while evaluation is performed in six zero-shot directions. IWSLT is a relatively small dataset, which tests our method’s ability to generalize from few languages.
@@ Appendix B Details of Our Experiments
Table 6: Summary of the language codes.
Europarl is a multilingual dataset crawled from the proceedings of the European Parliament. We again followed previous work (Liu et al., 2021) and evaluated our methods with a standard split for the zero-shot setting, containing English-centric data for eight languages with a total of 56 zero-shot evaluation directions. We specifically adopted the non-overlapping setting: in the original corpus, a sentence may be translated into multiple languages, and the non-overlapping setup chooses only one target translation for each input. This prevents potential data leaking problems. Europarl contains more data and languages than IWSLT, which further tests our method’s ability to generalize across multiple languages.
@@
In our presentation of beam search and the proposed EBBS, we describe the scorer as the multiplication of step-wise probabilities. In implementation, we used the sum of log-probabilities for numerical stability. Moreover, our EBBS is built on top of the popular fairseq framework (Ott et al., 2019), using their beam search implementation as the backbone. Consequently, we inherit standard beam search implementation techniques such as length normalization and max length constraints, which are not detailed in our pseudocode.
Beam Size BLEU BLEU1 BLEU2 BLEU3 BLEU4
1 24.98 56.75 30.08 18.84 12.46
3 25.99 57.04 30.92 19.65 13.17
[…]
@@
Appendix C Analysis of Beam Size
We analyze the effect of the beam size and show the results in Table 7. As seen, performance rises with a larger beam size and eventually plateaus at around five, which is consistent with the practice of standard beam search Meister et al. (2020). This analysis also shows the robustness of our EBBS algorithm. Based on this experiment and efficiency considerations, we used a beam size of five in our main experiments.
It should be mentioned that in principle our lower-level and upper-level beams need not have the same size. In our experiments, we simply set them to be the same. We have not tried full-grid hyperparameter tuning because the focus of our paper is algorithmic design, and as shown, our proposed EBBS is mostly robust to the beam size.
Figure 3: Inference time analysis on the test set of Italian to Dutch translation from the Europarl dataset. Experiments were conducted on an AMD EPYC 7313 CPU and an NVIDIA RTX A6000 GPU, with an inference batch size of 300 samples.
Appendix D Analysis of Inference Efficiency
We analyze the efficiency of our ensemble approach. As seen in Figure 3, the inference scales almost linearly, which is reasonable as we need to perform inference for all the components. The trend shows that it is computationally feasible to build an ensemble of even more components.
Further, the analysis suggests that our EBB-distilled models achieves a speedup of multiple times depending on the number of ensemble components. Recall from §4.3 that our our EBBS-based distillation is able to transfer the ensemble’s knowledge to a single multilingual model, without hurting, if not improving, the translation quality.
@@ Appendix D Analysis of Inference Efficiency
Appendix E Entropy of Distilled Models
We would like to understand why EBBS-based distillation largely outperforms other methods, such as union distillation (§4.3). The hypothesis is that cross-entropy distillation loss with diverse samples may lead to an overly smooth distribution, which in turn would affect the model performance Wen et al. (2023b); Shayegh et al. (2024).
Here, a large entropy generally indicates that the model is less certain, producing a more uniform prediction, whereas a low entropy indicates that the model is confident, producing a more peaked distribution.
Union distillation trains the model from the union of ensemble components’ outputs. It reduces prediction entropy compared with no distillation, but due to the nature of cross-entropy loss, it remains the highest among all distillation variants. Self-distillation is based on direction translation only, reinforcing the model’s current belief and thus producing the lowest entropy. On the contrary, our EBBS-based distillation achieves a moderate entropy on both datasets.
It should be emphasized that the entropy analysis only merely light on how different distillation methods behave, but the entropy itself does not indicate the quality of a model. We quote BLEU scores from Table 5, which has suggested that our EBBS-based distillation achieves similar or higher performance compared with EBBS, consistently outperforming other distillation methods.
Appendix F Case Study

### v2

@@ 5 Conclusion
In this work, we address ensemble-based zero-shot machine translation by directly translating and pivoting through different languages. We further design a novel bi-level beam search algorithm (called EBBS) for decoding. We evaluated EBBS on two popular zero-shot translation datasets, IWSLT and Europarl. Results show that EBBS outperforms existing ensemble techniques, and that the high-quality translations produced by EBBS can be used for distillation to improve translation efficiency (and sometimes also output quality).
Acknowledgments
The research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), a Mitacs Accelerate project, the Amii Fellow Program, the Canada CIFAR AI Chair Program, an Alberta Innovates Program, and the Digital Research Alliance of Canada (alliancecan.ca). We used the open-domain artwork “The Great Wave off Kanagawa” by Katsushika Hokusai in the title and as the background for Figure 1.
@@ 12       $\triangleright$ check for termination
Appendix B Experimental Details
Dataset details. We evaluated our methods using IWSLT 2017 (Cettolo et al. 2017) and Europarl v7 (Koehn 2005). Table 5 provides a summary of the languages.
The IWSLT 2017 translation dataset features multilingual data derived from TED talks. We followed previous work and used a standard split for zero-shot evaluation (Dabre and Kurohashi 2017; Liu et al. 2021). In particular, IWSLT contains English-centric training data for Italian, Dutch, and Romanian, while evaluation is performed in six zero-shot directions. IWSLT is a relatively small dataset, which tests our method’s ability to generalize from few languages.
@@ Appendix B Experimental Details
Table 5: The languages in the IWSLT and Europarl datasets.
Europarl is a multilingual dataset crawled from the proceedings of the European Parliament. We again followed previous work (Liu et al. 2021) and evaluated our methods with a standard split for the zero-shot setting, containing English-centric data for eight languages with a total of 56 zero-shot evaluation directions. We adopted their non-overlapping setting: in the original corpus, a sentence may be translated into multiple languages, and the non-overlapping setup chooses only one target translation for each input. This prevents potential data-leaking problems. Europarl contains more data and languages than IWSLT, which further tests our method’s ability to generalize across multiple languages.
@@
In our presentation of beam search and the proposed EBBS, we describe the scorer as the multiplication of step-wise probabilities. In implementation, we used the sum of log-probabilities for numerical stability. Moreover, our EBBS is built on top of the popular fairseq framework (Ott et al. 2019), using their beam search implementation as the backbone. Consequently, we inherit standard beam search implementation techniques such as length normalization and max length constraints, which are not detailed in our pseudocode.
Figure 3: Inference time analysis on the test set of Italian-to-Dutch translation from Europarl. Experiments were conducted on an AMD EPYC 7313 CPU and an NVIDIA RTX A6000 GPU, with a batch size of 300 samples.
Dataset Model BLEU chrF2++ TER COMET
IWSLT Direct translation (our replication) 17.29 42.57 72.46 0.7242
Pivoting (en) 16.19 42.79 80.39 0.7184
Word-level averaging ensemble 17.28 42.63 72.66 0.7237
Word-level voting ensemble 16.99 42.55 75.23 0.7162
Sequence-level voting ensemble (MBR) 16.72 42.50 76.12 0.7170
EBBS (ours) 18.24 43.66 73.50 0.7383
Europarl Direct translation (our replication) 27.85 54.37 63.40 0.8426
Pivoting (en) 27.75 54.08 63.03 0.8480
Word-level averaging ensemble 27.78 54.45 63.80 0.8409
Word-level voting ensemble 27.62 54.05 63.11 0.8409
Sequence-level voting ensemble (MBR) 28.14 54.43 62.75 0.8452
EBBS (ours) 28.44 54.64 62.48 0.8488
Table 6: Average performance statistics across all zero-shot language pairs in IWSLT and Europarl.
Lower Upper BLEU BLEU1 BLEU2 BLEU3 BLEU4 chrF2++ TER COMET
1 1 24.98 56.75 30.08 18.84 12.46 51.97 66.13 0.8255
3 3 25.99 57.04 30.92 19.65 13.17 52.65 65.60 0.8326
5 5 26.10 57.07 31.00 19.76 13.28 52.75 65.63 0.8340
7 7 26.10 57.06 30.98 19.74 13.30 52.75 65.69 0.8346
9 9 26.12 57.02 30.99 19.78 13.31 52.79 65.76 0.8352
5 1 25.06 56.70 30.10 18.88 12.49 52.07 66.20 0.8264
5 2 25.63 56.90 30.56 19.29 12.87 52.48 65.92 0.8311
5 3 25.94 56.93 30.87 19.61 13.14 52.69 65.79 0.8330
5 4 26.03 57.02 30.95 19.69 13.21 52.74 65.75 0.8332
5 5 26.10 57.07 31.00 19.76 13.28 52.75 65.63 0.8340
@@
Appendix C Analysis of Inference Efficiency
We analyze the efficiency of our ensemble approach. As seen in Figure 3, the inference scales almost linearly, which is reasonable as we need to perform inference for all the components. The trend shows that it is computationally feasible to build an ensemble of even more components. Importantly, recall this linearly increasing inference time is mitigated by EBBS distillation (Row 8, Table 1), which is as fast as one component.
Appendix D Average Performance across Tasks
In addition to the main results (Table 1), we further show the average performance across the 6 and 56 translation directions on IWSLT and Europarl using various metrics. We omit EBBS distillation, because distilling across all directions requires significant computing resources. As shown in Table 6, EBBS consistently outperforms existing baselines, with the only exception for TER on the IWSLT dataset, where direct translation achieved a higher performance. This is potentially because the TER score is based on edit distance, which is not robust to word reordering. In general, results show EBBS consistently outperforms competing methods.
Appendix E Analysis of Beam Size
We analyze the effect of different beam sizes on our EBBS algorithm. First, we study the setting where the lower- and upper-level beam sizes are matched. As seen in the top half of Table 7, the performance tends to increase with a larger beam size and eventually plateaus at around five, which is consistent with the practice of standard beam search (Meister, Cotterell, and Vieira 2020).
Further, we analyze the setting where the upper- and lower-level beam sizes are not matched. Generally, the upper-level beam size should not exceed the lower-level beam size, because otherwise the upper-level beam may not be fully filled by the ensemble components. As shown in the bottom half of Table 7, EBBS performs better with larger upper-level beam sizes. This is understandable because a larger upper-level beam allows EBBS to explore more candidates in general.
Overall, our analysis shows that EBBS is robust and works well with a variety of beam sizes. Based on this experiment and efficiency considerations, we used a beam size of five for both upper- and lower-level beams in our main experiments.
@@ Appendix E Analysis of Beam Size
Voting scheme BLEU BLEU1 BLEU2 BLEU3 BLEU4 chrF2++ TER COMET
None (beam search) 25.33 56.32 30.08 19.01 12.78 52.32 66.56 0.8276
Total-sum 25.27 56.67 30.27 19.07 12.65 52.19 66.12 0.8311
Max 25.81 56.89 30.76 19.51 13.09 52.46 65.92 0.8300
0/1 25.84 56.99 30.78 19.49 13.05 52.61 65.80 0.8322
Top-$Z$ sum (ours) 26.10 57.07 31.00 19.76 13.28 52.75 65.63 0.8340
Table 9: Comparison of different ensemble variants, using Italian-to-Dutch translation in the Europarl dataset as the testbed.
Appendix F Entropy of Distilled Models
We would like to understand why EBBS-based distillation largely outperforms other methods, such as union distillation (§4.3). Our hypothesis is that cross-entropy distillation loss with diverse samples may lead to an overly smooth distribution, which in turn would affect the model performance (Wen et al. 2023b; Shayegh et al. 2024).
A large entropy generally indicates that the model is less certain, producing a more uniform prediction, whereas a low entropy indicates that the model is confident, producing a more peaked distribution.
Union distillation trains the model from the union of ensemble components’ outputs. It reduces prediction entropy compared with no distillation, but due to the nature of cross-entropy loss, it remains the highest among all distillation variants. Direct distillation is based on direct translation only, reinforcing the model’s current belief and thus producing the lowest entropy. On the contrary, our EBBS-based distillation achieves a moderate entropy on both datasets.
It should be emphasized that the entropy analysis merely shed light on how different distillation methods behave, but the entropy itself does not indicate the quality of a model. We quote BLEU scores from Table 3, which has suggested that our EBBS-based distillation achieves similar or higher performance compared with EBBS, consistently outperforming other distillation methods.
Appendix G Analysis of Voting Methods in EBBS
In our EBBS algorithm, the lower-level beams are synchronized into a shared upper-level beam by voting. Specifically, EBBS uses a mechanism of top-$Z$ sum voting, where we add the ensemble components’ probabilities for each appearance of a candidate in the lower-level beam, shown in Eqn. (3). Here, we analyze a few alternative voting methods for EBBS.
If EBBS adopts total-sum voting, it still uses lower-level beams to find candidates, but adds all components’ probabilities together. This is equivalent to applying the common averaging ensemble to the top-$Z$ candidates. However, it differs from our approach, because in total-sum voting, a component will vote even if the candidate does not appear in its own lower-level beam; the probability after voting in Eqn. (3) is substituted with $\frac{1}{K}\sum_{k}p_{k}(\mathbf{y}|\mathbf{x})$. As shown in Table 9, EBBS with total-sum voting performs worse than direct translation, suggesting the importance of ignoring the components whose lower-level beam does not contain the candidate. This is analogous to nucleus sampling (Holtzman et al. 2019), where the long tail of a distribution is mainly noise and should be ignored.
Other voting schemes that EBBS may use include 0/1 voting and max voting. The former selects the candidates that appear most in the lower-level beams, disregarding the probability values (unless for ties); the latter chooses the maximum probability across the lower-level beams, which gives preference to sequences through a maximization bias (Hasselt 2010; van Hasselt, Guez, and Silver 2016). As seen, EBBS performs relatively well with both of these voting schemes, achieving a decent improvement over the baseline approach; however, their performance is worse than our top-$Z$ sum voting.
Overall, the proposed bi-level beam search ensemble is effective with different voting schemes (except for the total-sum voting), and our top-$Z$ sum voting works the best among these variants.
@@ Appendix G Analysis of Voting Methods in EBBS
Appendix H Case Study